8. Liquidations in bankruptcy Aa Aa When a business is worth more if its assets are sold than if it continues to operate, then the business is liquidated, and the proceeds from the sale are used to satisfy any outstanding debt. Liquidation occurs when businesses file for bankruptcy under Chapter 7 of the Federal Bankruptcy Reform Act. This act provides for an equitable distribution of the debtor's assets among the creditors. The distribution of assets is governed by a certain priority of claims. Which of the following claimants has the highest priority according to Chapter 7? Unfunded pension plan liabilities Preferred stockholders O Unsecured claims for customer deposits Common stockholders Celestial Crane Cosmetics has been in financial distress for the past three years. The company's reorganization plans rejected, and the bankruptcy court is forcing Celestial Crane Cosmetics to liquidate its assets and settle creditor claims were The court appointed the trustee with an estimated expense budget of $281,250. Subordinated debentures are subordinated only to the notes payable. Celestial Crane Cosmetics Balance Sheet (Thousands of dollars) $1,875 Accounts payable $375 Current assets Net fixed assets Notes payable 1,875 750 Subordinated debentures 750 Total debt 1,875 Common equity 1,875 Total liabilities and equity Total assets $3,750 $3,750 The trustee has arranged the sale of Celestial Crane Cosmetics's assets and estimates to receive $2.50 million in proceeds from the liquidation.
Proceeds from the liquidation will be used to settle Celestial Crane Cosmetics's liabilities. Based on your understanding of distribution of assets to various claimants, how much will common stockholders receive (if they receive anything at all) when Celestial Crane Cosmetics is liquidated? $718,750 $625,000 $378,125 $343,750 The bankruptcy courts and the company have discretion in how they elect to handle a failing firm. Just as the decision to reorganize or liquidate the firm depends on the relationship between its going concern and its liquidation values, the significance and competitiveness of the industry in which the failing firm operates also affects the court's decision. If Celestial Crane Cosmetics were a public utility company or a railroad service company instead of an industrial corporation, the bankruptcy courts would likely push for: Liquidation Reorganization
Answer #1

1). Unfunded pension liabilities will have the first claim since priority claims have to be settled first. Unsecured claims are settled only after that.

2). Common stockholders will receive $343,750.

Proceeds from sale of assets = 2,500,000

Net proceeds after deducting expenses = 2,500,000 - 281,250 = 2,218,750

A/C payable, notes payable and subordinated debt will all be paid off first before common stockholders.

So, amount left for common stockholders = 2,218,750 - 1,875,000 = 343,750

3). The courts will most likely push for reorganization (under Chapter 11) as public companies can control the reorganization process and still run their business, so that employees and other creditors are minimally affected.